The market value of ABC company's common stock is $20 million. It has $5 million in debt with an interest rate of 6.25%. The beta of the company's common stock is 1.25, and the market risk premium is 8%. If the Treasury bill rate is 5% and the company’s tax rate is 20%, what is the company's cost of capital?

a. 13.0 percent b. 14.6 percent c. 15.0 percent d. 7.0 percent

The historical returns for the past three years for Stock B and the stock market were the following: Stock B: 2016: 24%, 2017: 0%, 2018: 24%; market return: 2016: 10%, 2017: 12%, 2018: 20%. Calculate the beta for Stock B.

a. 0.86 b. 1.00 c. 1.17 d. 1.13

Company A's historical returns for the past three years were 6%, 15%, and 15%. Similarly, the market return were 10%, 10%, and 16%. Suppose the risk-free rate of return is 4% and that investors expect the market to return 10%. What is the company’s cost of equity computed with the CAPM?

a. 14.0 percent b. 8.5 percent c. 12.0 percent d. 10.0 percent

A company’s capital structure is 4 million Euros in debt and 16 million Euros in equity

If the beta of its debt is 0.25 and the beta of its equity is 1.50, what is the company’s asset beta

a. 0.75 b. 1.25 c. 1.10 d. 1.19

The market value of DEF Company's equity is $15 million and the market value of its debt is $5 million. The required rate of return on its equity is 20% and 8% on its debt, Calculate the company's cost of capital if its tax rate is 25%

a. 20 percent b. 16.5 percent c. 14.5 percent d. 17 percent
Answer #1

O given a debt = $5 billion > Wa= 5= 951- 20 & we = 100% - 25%. = ord 75-). Ca= 6.25-1 (1-20..) = 50% Caffer tax) Ce = Solet

I can only answer 1 question at a time so I am solving am solving question 1. Please do rate me and mention doubts in the comments section.